THE RANDOLPHS USED 12 MORE GALLONS OF FUEL OIL IN OCTOBER THAN IN SEPTEMBER AND TWICE AS MUCH OIL IN NOVEMBER AS IN SEPTEMBER. IF THEY USED 132 GALLONS FOR THE THREE MONTHS, HOW MUCH WAS USED DURING EACH MONTH?
THANKS FOR ANY HELP.

Answer by funmath(2933)

THE RANDOLPHS USED 12 MORE GALLONS OF FUEL OIL IN OCTOBER THAN IN SEPTEMBER AND TWICE AS MUCH OIL IN NOVEMBER AS IN SEPTEMBER. IF THEY USED 132 GALLONS FOR THE THREE MONTHS, HOW MUCH WAS USED DURING EACH MONTH?
:
Let the amount used in Sept.=x
Then the amount used in Oct.=x+12
Then the amount used in Nov.=2x
Add the months together=132
:
Equation to solve:
(x)+(x+12)+(2x)=132
4x+12=132
4x+12-12=132-12
4x=120
4x/4=120/4
x=30
:
Sept: x=30 gal
Oct:x+12=30+12=42 gal
Nov:2x=2(30)=60 gal
:
Check, if you add all three months do you get 132 gal?
30+42+60=132
132=132 Yes it does!!!
Happy Calculating!!!